you are viewing a single comment's thread
view the rest of the comments
[–] -5 points 2 months ago* (last edited 2 months ago) (4 children)

Chance == odds

so

4/4 == totalChoice_permitted 
one_Choice = 1/4 

A=25
B=50
C=0
D=25

if one_Choice  ==  one_Choice
then combine 
	reduce choices by the amount simlair plus one

	A==D
	combine

3/3 == totalChoice_permitted 
one_Choice = 1/3 
	turn to decimal (one_Choice) x 100 == 33%
	

1 = AD =25%
2 = B =50%
3 = C =0%

if answer not equal to (one_Choice)
	print “the answer is not listed but the answer is in the struggle of answering a question so 0%”




Not C but 0% Not all odds matter or they do. It’s a double whammy

or do I got pie on my face?

Is it the percentage or the place holder?

What are the rules?

I would think it would be safe to assume the limits are out of four choices but when two are the same.

like linear or some shit.

Update: I asked my local AI clown clanker and it gave this as a answer

A clever question! Since the answer is chosen at random, each option has an equal chance of being correct. Therefore, the probability that I will be correct if I pick randomly is: 1/4 = 25% So, the correct answer is (a) 25%!

The olgipity is very sure of A) and not D)

  • source
  • hideshow 4 child comments
  • [–] 4 points 2 months ago (2 children)

    So if you are playing Russian roulette with 5 loaded chambers, you can ignore 4 of them because the outcome is the same as one of the other chambers, meaning you have a 50/50 chance of winning. Truly astounding math.

  • source
  • parent
  • hideshow 2 child comments